# Classifying representations of finite classical groups of Lie type of dimension up to $\ell^{4}$

### v1

@@ 00 $1$
Suppose that $\ell\geq 12$. The dimensions of the weight spaces associated to the highest weights 00, $\lambda_{\ell}$, and $2\lambda_{\ell}$, are computed in Table 2 of [11]. We next shall compute the dimension of $V(3\lambda_{\ell})$, the dimension of $V(4\lambda_{\ell})$ follows by means of the same computations. To this end and according to Proposition 2.1 we have to compute $l(\mu)$, for each dominant weight $\mu\preceq 3\lambda_{\ell}$ given in Lemma 5.3. We first observe that the stabilizers of the weights $\mu_{i}=i\lambda_{\ell}$, for $i=1,2,3$, are reflection groups of type $B_{\ell-1}$, and hence $l(\mu_{i})=2\ell$. On the other hand, we also see that the stabilizers of $\lambda_{\ell-1}+\lambda_{\ell}$, $\lambda_{\ell-1}$ and $\lambda_{\ell-2}$ are reflection groups of type $B_{\ell-2}$, $B_{\ell-2}\times A_{1}$ and $B_{\ell-3}\times A_{2}$, so the length of its orbits are $4\ell(\ell-1)$, $2\ell(\ell-1)$, $34\ell(\ell-1)(\ell-2)/$ 3 respectively and clearly $l(0)=1$. Then, by resorting to the multiplicities given in Lemma 5.3 and Proposition 2.1 we get
$\dim V(3\lambda_{\ell})$ $=2\ell+4\ell(\ell-1)+4\ell(\ell-1)(\ell-2)/3+2\ell+2\ell(\ell-1)+2\ell^{2}+\ell$
$=\ell(2\ell+1)(2\ell+5)/3.$
Now we compute the dimension of $V(\lambda_{\ell-3})$ since the case of $V(\lambda_{\ell-2})$ follows in the same way, even shorter, and $V(\lambda_{\ell-1})$ was computed in [11]. Certainly $\dim V(\lambda_{\ell-3})$ may be calculated as above, but let us give an alternative way to deduced it. We first observe that
$\dim V(\lambda_{\ell-3})\geq m_{\lambda_{\ell-3}}(\lambda_{\ell-3})l(\lambda_{\ell-3})+m_{\lambda_{\ell-3}}(\lambda_{\ell-2})l(\lambda_{\ell-2})+m_{\lambda_{\ell-3}}(\lambda_{\ell-1})l(\lambda_{\ell-1})+m_{\lambda_{\ell-3}}(\lambda_{\ell})l(\lambda_{\ell})+m_{\lambda_{\ell-3}}(0)l(0).$
As above we see that the stabilizer of $\lambda_{\ell-i}$ is a reflection group of type $B_{\ell-i-1}\times A_{i}$, for $i=1,2,3$ and of type $B_{\ell-2}$ for $\lambda_{\ell}$. Then, $l(\mu)$ is 2(1)(2)(3)/3, 4(1)(2)/3, 2(1), 2, $12\ell(\ell-1)(\ell-2)(\ell-3)/3,\>4\ell(\ell-1)(\ell-2)/3,\;2\ell(\ell-1),\;2\ell,\;$ 1 for $\mu=\lambda_{\ell-3},\;\mu=\lambda_{\ell-2},\;\mu=\lambda_{\ell-1},\;\mu=\lambda_{\ell},\;\mu=0$, respectively. Thus, by using the multiplicities given by Lemma 5.3 we see that
$\dim V(\lambda_{\ell-3})$ $\geq 2\ell(\ell-1)(\ell-2)(\ell-3)/3+4\ell(\ell-1)(\ell-2)/3+2\ell(\ell-1)(\ell-2)+2\ell(\ell-1)+\ell(\ell-1)/2$
$=(2\ell-2)(2\ell-1)2\ell(2\ell+1)$
$=\dim\Lambda^{4}(V).$
By combining the last inequality with $V(\lambda_{\ell-3})\subseteq\Lambda^{4}(V)$ we conclude that $V(\lambda_{\ell-3})$ is equal to $\Lambda^{4}(V)$.
On the other hand, we can deduce the dimension of $V(\lambda_{\ell-1}+\lambda_{\ell})$ by appealing Lemma 5.5 and arguing as in $V(3\lambda_{\ell})$.
Finally, for $\lambda=\lambda_{1}$ and $12\leq\ell\leq 15$ we infer that there not exist $\mu\preceq\lambda$ except $\lambda$ itself. In addition, we directly verify that $\rm{Orb}_{\mathcal{W}}(\lambda)=2^{\ell}$. This then says that $\dim V(\lambda)=2^{\ell}<\ell^{4}$ for $\ell$ in the mentioned rank, i.e., $\lambda$ is admissible. ∎
Remark 5.7.
Let $G$ be as above of type $B_{\ell}$ and $V$ be the irreducible $G$-module of dimension $12\ell+$ 1. Then, the anti-symmetric tensor product $\Lambda^{i}(V)$ is irreducible, for $i=1,2,3,4$.
In order to prove that the weights in the previous proposition are the only admissible weights of type $B_{\ell}$, we will prove the following combinatorial lemma which will be very useful below. This result is perhaps well-known to the experts, but we give a proof for lack of a reference.
@@
Firstly we put $t_{k}:=2^{n-k}\binom{n}{k}$ for each integer $0\leq k\leq n$. Then
@@
We begin by observing $t_{0}=2^{n-1}\cdot 2<2^{n-1}\cdot n=t_{1}$ because $2<n$. We also note that
@@
since $22<(n-1)/$ 2. Thus, $t_{0}<t_{1}<t_{2}$. In general, if $k<\lfloor\frac{n}{3}\rfloor$ we may infer that $2k<(n-(k+1))$. It works for $n\geq 12$ and proceed as above to show $t_{k}<t_{k+1}$. We see that the worst case is when $k=\lfloor\frac{n}{3}\rfloor-1$. In such case, the inequality follows from
$n-\left(\left\lfloor\frac{n}{3}\right\rfloor-1\right)\geq n-\frac{n}{3}+1=2\frac{n}{3}+1\geq 2\left\lfloor\frac{n}{3}\right\rfloor.$
We then deduce
@@
$=2^{n-\lfloor\frac{n}{3}\rfloor+1}\frac{n!}{(\lfloor\frac{n}{3}\rfloor-1)!(n-\lfloor\frac{n}{3}\rfloor+1)!}$
@@
thanks to the inequality $2/(n-\lfloor\frac{n}{3}\rfloor+1)<1/\lfloor\frac{n}{3}\rfloor$. This proves that the coefficients of $(2+x)^{n}$ are strictly increasing until $k=\lfloor\frac{n}{3}\rfloor$.
On the other hand, by similar arguments, when $k>\lfloor\frac{n}{3}\rfloor$ we have that $2\geq\frac{n-(k-1)}{k}$. Then, by observing,
@@
Now we are ready to state and give a proof of our main result of this section.
@@
Let $G$ be a finite group of Lie type $B_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 12$ then the list giving in Proposition 5.6 exhausts the admissible weights.
@@
Nullity of coefficients -- $\lambda$ is admissible implies $a_{i}=0$ for $41\leq i\leq\ell-$ 4.
Suppose first that $a_{i}\neq 0$, for some $42\leq i\leq\ell-$ 4. Then, the stabilizer of $\lambda$ is contained in a reflection group of type $B_{i-1}\times A_{\ell-i}$ it is worth mentioning that we are using the convention $B_{1}=A_{1}$ as in [6]. This implies
$\dim V(\lambda)\geq t_{i}:=\frac{2^{\ell}\ell!}{2^{i-1}(i-1)!(\ell-i+1)!}$ $=2^{\ell-(i-1)}\binom{\ell}{i-1}.$
By appealing to Lemma 5.8 we have that the $(i-1)$th-coefficient $t_{i}$ of $(2+x)^{\ell}$ are strictly increasing for $12\leq i\leq\lfloor\frac{\ell}{3}\rfloor+$ 1. Thus, for $\ell\geq 16$ we use the fact that $2^{\ell}\geq\ell^{4}$ to deduce
$\dim(V(\lambda))\geq t_{i}>t_{0}=2^{\ell}\geq\ell^{4}.$
Now, by numerical inspection we also may verify for each $12\leq\ell\leq 15$ that
$\dim V(\lambda)\geq$ $1\displaystyle 2^{\ell-(i-1)}\binom{\ell}{i-1}>\ell^{4},\quad\quad\text{ for }2\leq i\leq\lfloor\frac{\ell}{3}\rfloor+$ 1 .
On the other hand, we will demonstrate that $\dim V(\lambda)>\ell^{4}$ by using the sequence $(t_{i})$ is decreasing for $i>\lfloor\frac{\ell}{3}\rfloor+1$. Indeed, we shall show that $t_{\ell-4}>\ell^{4}$ for $\ell\geq 12$ by observing
@@
Since $\ell\geq 12$ we can infer that $121-j/\ell\geq(12-j)/$ 12 for $1\leq j\leq 3$, and $(\ell-4)/15\geq 8/15$. By combining these facts we can infer
@@
$\dim(V(\lambda))\geq\frac{|\mathcal{W}|}{|\rm{Stab}_{\mathcal{W}}(\lambda^{\prime})|}>\ell^{4},$
i.e., $\lambda$ is not admissible when $a_{1}\geq 2$.
By appealing to the same arguments, $a_{1}=1$ implies $a_{i}=0$ for any $2\leq i\leq\ell$. Hence we reduce to
@@
By Proposition 5.6 we have that $\lambda=\lambda_{1}$ is admissible when $12\leq\ell\leq 15$ and not otherwise. Next we consider $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. If $a_{\ell-3}\geq 2$, we produce a dominant weight $\lambda^{\prime}:=\lambda-\alpha_{\ell-3}$, whose writing as nonnegative linear combination of fundamental weights contains the term $\lambda_{\ell-4}$, but it cannot be possible due to the arguments above. Thus, we reduce the analysis to
@@
$\dim V(\lambda)\geq\frac{|\mathcal{W}|}{|\rm{Stab}_{\mathcal{W}}(\lambda)|}>\ell^{4}.$
@@
$\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}\quad\text{ or }\quad\lambda=a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}.$
In both cases we may deduce that $\lambda$ is not admissible if $a_{\ell-i}>1$ with $i=0$ or $1$ by arguing as above. For instance, if $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. with $a_{\ell-1}\geq 2$ we can observe that the term $2\lambda_{\ell-2}$ appears in $\lambda^{\prime}:=\lambda-\alpha_{\ell-1}$, but such a weight $\lambda^{\prime}$ is not admissible by the previous case. It follows that $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. Moreover, we may proceed very similarly for $a_{\ell}$ which implies to analyze the dimensions of the representations associated to the weights
$\lambda_{\ell-2}+\lambda_{\ell-1},\;\lambda_{\ell-2}+\lambda_{\ell}\;\text{ and }\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$
@@
we deduce that $\mu:=\lambda_{\ell-3}+\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}$. The same calculations as above show that the length of orbit of $\mu$ is greater to $\ell^{4}$, and hence $\dim V(\lambda)>\ell^{4}$, i.e., $\lambda$ is not admissible.
On the other hand, $\lambda:=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by Example 3.4. While for $\lambda:=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ one notices that
@@
which says $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda$. Thus, we have
$\dim V(\lambda)$ $\geq l(\lambda_{\ell-3}+2\lambda_{\ell})$
$=2^{3}\ell(\ell-1)(\ell-2)(\ell-3)/3$
$=\ell^{4}\cdot 2^{3}(1-1/\ell)(1-2/\ell)(1-3/\ell)/3$
$>\ell^{4},$
by observing $12^{3}(1-1/\ell)(1-2/\ell)(1-3/\ell)/3>$ 1 for any $\ell\geq 12$. This implies that $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ is not admissible either.
@@
forcing then $\dim V(\lambda)\geq l(\lambda_{\ell-4})>\ell^{4}$.
Therefore, the Table in Proposition 5.6 exhausts the admissible weights which conclude the proof. ∎
Finally, we will deal with the rest of the classic groups Lie type in the next subsection.
@@ 5.2. Type $C_{\ell}$ and $D_{\ell}$
Let $G$ be a group of type $C_{\ell}$ or $D_{\ell}$ over $\mathbb{\overline{F}}_{p}$. If $\ell\geq 12$ then the lists giving in the table here below exhaust the admissible weights in each case.
type $\lambda$ dimension
$C_{\ell}$
0 1
$\lambda_{\ell}$ $2\ell$
$2\lambda_{\ell}$ $(2\ell+1)(2\ell)/2$
$3\lambda_{\ell}$ $(2\ell+2)(2\ell+1)(2\ell)/6$
$4\lambda_{\ell}$ $(2\ell+3)(2\ell+2)(2\ell+1)(2\ell)/24$
$\lambda_{\ell-1}$ $12\ell^{2}-\ell-$ 1
$\lambda_{\ell-2}$ $2\ell(2\ell-1)(2\ell-2)/6-2\ell$
$\lambda_{\ell-3}$ $((2\ell-3)(2\ell-2)(2\ell-1)2\ell/24)-((2\ell+1)(2\ell)/2)$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+2)(2\ell)(2\ell-2)/3$
$D_{\ell}$
0 1
$\lambda_{\ell}$ $2\ell$
$2\lambda_{\ell}$ $(2\ell+1)(2\ell)/2-1$
$3\lambda_{\ell}$ $(2\ell+2)(2\ell+1)(2\ell)/6-2\ell$
$4\lambda_{\ell}$ $((2\ell+3)(2\ell+2)(2\ell+1)(2\ell)/24)-((2\ell+1)(2\ell)/2)$
$\lambda_{\ell-1}$ $2\ell^{2}-\ell$
$\lambda_{\ell-2}$ $62\ell(2\ell-1)(2\ell-2)/$ 6
$\lambda_{\ell-3}$ $(2\ell-3)(2\ell-2)(2\ell-1)2\ell/24$
$\lambda_{\ell}+\lambda_{\ell-1}$ $(2\ell+2)(2\ell)(2\ell-2)/3$
In addition, when $12\leq\ell\leq 17$ and $G$ is of type $D_{\ell}$, the second table has to include the weights $\lambda_{1}$ and $\lambda_{2}$ whose associated irreducible representation has dimension $2^{\ell-1}$.
@@
The tables in Appendix contain all the information to calculate directly the dimensions of the weights spaces $V(\lambda)$. Thus, we may deduce that the weights in the tables above are effectively admissible. Next we shall prove that the list is exhausted respect to admissible ones. To this end, let
@@
Contrary to the previous case, here, the weights $\lambda=\lambda_{1}$, $\lambda=\lambda_{2}$ are admissible for $12\leq\ell\leq 17$ since the dimension of $V(\lambda_{i})$ is $2^{\ell-1}$ for $i=1,2$. On the other hand, the dominant weights of the form $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ can be addressed as above by using Appendix -Table 2. ∎
Remark 5.11.
It is worth mentioning that:
  • If $G$ is of type $C_{\ell}$ and $V$ the irreducible $G$-module of dimension $2\ell$. Then, the $i$-symmetric power $Sym^{i}(V)$ is irreducible and the anti-symmetric tensor product $\Lambda^{i}(V)$ is reducible for $i=2,3,4$.
  • In contrast with the case $C_{\ell}$, when $G$ is of type $D_{\ell}$ we may observe that $\operatorname{Sym}^{i}(V)$ is reducible and $\Lambda^{i}(V)$ is irreducible for $i=2,3,4$.
6. Some applications
Let $\rho_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{Q}}_{p})\subseteq\operatorname{GL}_{n}(\overline{\mathbb{Q}}_{p})$ be a $p$-adic orthogonal Galois representation. Examples of this kind of Galois representations can be obtained from the so-called regular algebraic, polarizable, cuspidal automorphic representations of $\operatorname{GL}_{n}(\mathbb{A}_{\mathbb{Q}})$, via the global Langlands correspondence [2, §2.1].
In this case, the large image conjecture predicts that, for almost all primes $p$, the image of the proyectivizations $\overline{\rho}^{proj}_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{PGO}_{n}(\overline{\mathbb{F}}_{p})$ of the residual representations $\overline{\rho}_{p}:G_{\mathbb{Q}}\rightarrow\operatorname{GO}_{n}(\overline{\mathbb{F}}_{p})$ should contains $\operatorname{P\Omega}_{n}^{\pm}(\mathbb{F}_{p^{s}})$ for some positive integer $s$. As we pointed out in the introduction, this conjecture is known in very few cases. However, if we impose certain conditions on the local components of $\pi$, at least we can bound the possible images of $\overline{\rho}_{p}$. For example, in [14, Proposition 3.4], the second author proved the following criterion concerning to the image of certain families of orthogonal Galois representations.

### v3

@@ 0 $\ell(\ell-1)/2$
Suppose that $\ell\geq 12$ and $p\neq 2$. The admissibility of the weights $3\lambda_{\ell}$ will be deduced by bounding the dimension of $L(3\lambda_{\ell})$. The case of $4\lambda_{\ell}$ will follow similarly. To this end, and according to Eq. 2.4, we have a bound $m_{L(\lambda)}(\mu)$ for each dominant weight $\mu\preceq 3\lambda_{\ell}$ given in Lemma 4.3. On the other hand, observe that the stabilizers of the weights $\mu_{i}=i\lambda_{\ell}$, for $i=1,2,3$, are reflection groups of type $B_{\ell-1}$, which implies $l(\mu_{i})=2\ell$. In addition,we also see that the stabilizers of $\lambda_{\ell-1}+\lambda_{\ell}$, $\lambda_{\ell-1}$ and $\lambda_{\ell-2}$ are reflection groups of type $B_{\ell-2}$, $B_{\ell-2}\times A_{1}$ and $B_{\ell-3}\times A_{2}$, so the length of its orbits are $4\ell(\ell-1)$, $2\ell(\ell-1)$, $4\ell(\ell-1)(\ell-2)/3$ respectively and clearly $l(0)=1$. Then, by resorting to the multiplicities given in Lemma 4.3 and Eq. (2.4) we get
$\dim L(3\lambda_{\ell})$ $\leq 2\ell+4\ell(\ell-1)+4\ell(\ell-1)(\ell-2)/3+2\ell+2\ell(\ell-1)+2\ell^{2}+\ell$
$=\ell(2\ell+1)(2\ell+5)/3,$
which implies $3\lambda_{\ell}$ is admissible. Similarly we can bound the dimensions of $V(\lambda_{\ell-2})$ and $V(\lambda_{\ell-3})$ by using Lemma 4.2 and Eq. (2.4), and whence bounds for $L(\lambda_{\ell-2})$ and $L(\lambda_{\ell-3})$. Now, the weight $\lambda_{\ell-1}+\lambda_{\ell}$ is admissible by appealing to [13, Table 3]. Finally, when $12\leq\ell\leq 16$, we infer that there not exist $\mu\preceq\lambda_{1}$ except $\lambda_{1}$ itself which implies that $\lambda_{1}$ is admissible since $l(\lambda_{1})=2^{\ell}\leq\ell^{4}$. ∎
In order to get a complete list of admissible weights of type $B_{\ell}$, we will prove the following combinatorial lemma which will be very useful below. This result is perhaps well-known to the experts, but we give a proof for lack of a reference.
@@
Firstly, we set $t_{k}:=2^{n-k}\binom{n}{k}$ for each integer $0\leq k\leq n$. Then
@@
We begin by observing $t_{0}=2^{n-1}\cdot 2<2^{n-1}\cdot n=t_{1}$ because $2<n$. Notice also that
@@
thus $t_{0}<t_{1}<t_{2}$. We may proceed as above to show $t_{k}<t_{k+1}$ when $k<\lfloor\frac{n}{3}\rfloor$. It suffices to verify the limit case $k=\lfloor\frac{n}{3}\rfloor-1$. In such a case, from the inequality follows from
$n-\left(\left\lfloor\frac{n}{3}\right\rfloor-1\right)\geq n-\frac{n}{3}+1=2\frac{n}{3}+1\geq 2\left\lfloor\frac{n}{3}\right\rfloor,$
we can deduce that
@@
thanks to the inequality $2/(n-\lfloor\frac{n}{3}\rfloor+1)<1/\lfloor\frac{n}{3}\rfloor$. This proves that the coefficients of $(2+x)^{n}$ are strictly increasing until $k=\lfloor\frac{n}{3}\rfloor$. On the other hand, by similar arguments, when $k>\lfloor\frac{n}{3}\rfloor$ we have that $2\geq\frac{n-(k-1)}{k}$. Then, by observing,
@@
Now we establish the main result concerning admissible weights of type $B_{\ell}$.
@@
Let $G$ be a finite group of Lie type $B_{\ell}$ over $\mathbb{\overline{F}}_{p}$ with $p>2$. If $\ell\geq 12$ then the following list of weights
$0,\lambda_{\ell},\;2\lambda_{\ell},\;3\lambda_{\ell},\;4\lambda_{\ell},\;\lambda_{\ell-1},\;\lambda_{\ell-2},\;\lambda_{\ell-3},\;\lambda_{\ell-1}+\lambda_{\ell}\mbox{ and }\lambda_{1}\;(\mbox{when }12\leq\ell\leq 16)$
exhausts the admissible weights for $B_{\ell}$.
@@
Nullity of coefficients - $\lambda$ is admissible implies $a_{i}=0$ for $1\leq i\leq\ell-4$.
Suppose first that $a_{i}\neq 0$, for some $2\leq i\leq\ell-4$. Then, the stabilizer of $\lambda$ is contained in a reflection group of type $B_{i-1}\times A_{\ell-i}$ - it is worth mentioning that we are using the convention $B_{1}=A_{1}$ as in [14]. This implies
$\dim L(\lambda)\geq t_{i}:=\frac{2^{\ell}\ell!}{2^{i-1}(i-1)!(\ell-i+1)!}$ $=2^{\ell-(i-1)}\binom{\ell}{i-1}$.
By numerical inspection we also may verify for each $12\leq\ell\leq 15$ that
$\dim L(\lambda)\geq$ $2^{\ell-(i-1)}\binom{\ell}{i-1}>\ell^{4},\quad\quad\text{ for }2\leq i\leq\lfloor\frac{\ell}{3}\rfloor+1.$
Now, for $\ell\geq 16$, we observe that
$\dim L(\lambda)\geq t_{i}>t_{0}=2^{\ell}\geq\ell^{4}.$
by using the (i1)(i-1)th-coefficients $t_{i}$ of $(2+x)^{\ell}$ are strictly increasing for $2\leq i\leq\lfloor\frac{\ell}{3}\rfloor+1$ and $2^{\ell}\geq\ell^{4}$.
On the other hand, we will demonstrate that $\dim L(\lambda)>\ell^{4}$ by using the sequence $(t_{i})$ is decreasing for $i>\lfloor\frac{\ell}{3}\rfloor+1$. Indeed, we shall show that $t_{\ell-4}>\ell^{4}$ for $\ell\geq 12$ by observing
@@
Since $\ell\geq 12$ we can infer that $1-j/\ell\geq(12-j)/12$ for $1\leq j\leq 3$, and $(\ell-4)/15\geq 8/15$. By combining these facts we can infer
@@
$\dim L(\lambda)\geq\frac{|\mathcal{W}|}{l(\lambda^{\prime})}>\ell^{4},$
i.e., $\lambda$ is not admissible when $a_{1}\geq 2$. By appealing to the same arguments, $a_{1}=1$ implies $a_{i}=0$ for any $2\leq i\leq\ell$. Hence we reduce to
@@
By Proposition 4.4 we have that $\lambda=\lambda_{1}$ is admissible when $12\leq\ell\leq 16$. On the other hand, since $\dim L(\lambda)=2^{\ell}>\ell^{4}$ for $\ell>16$ by [13, Table 8], $\lambda_{1}$ is not admissible. Next we consider $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. If $a_{\ell-3}\geq 2$, we produce a dominant weight $\lambda^{\prime}:=\lambda-\alpha_{\ell-3}$, whose writing as nonnegative linear combination of fundamental weights contains the term $\lambda_{\ell-4}$, but it cannot be possible due to the arguments above. Thus, we reduce the analysis to
@@
$\dim L(\lambda)\geq\frac{|\mathcal{W}|}{l(\lambda^{\prime})}>\ell^{4}.$
@@
[…]
In both cases we may deduce that $\lambda$ is not admissible if $a_{\ell-i}>1$ with $i=0$ or $1$ by arguing as above. For instance, if $\lambda=\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. with $a_{\ell-1}\geq 2$ we can observe that the term $2\lambda_{\ell-2}$ appears in $\lambda^{\prime}:=\lambda-\alpha_{\ell-1}$, but such a weight $\lambda^{\prime}$ is not admissible by the previous case. It follows that $\lambda=\lambda_{\ell-2}+a_{\ell}\lambda_{\ell}$ or $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$. In those cases we may proceed very similarly to rule out the weights when $a_{\ell}>1$. Thereby, we have to examine the admissibility of:
[…]
@@
we deduce that $\mu:=\lambda_{\ell-3}+\lambda_{\ell}\prec\lambda_{\ell-2}+\lambda_{\ell-1}$. The same calculations as above show that the length of orbit of $\mu$ is greater to $\ell^{4}$, and hence $\dim L(\lambda)>\ell^{4}$, i.e., $\lambda$ is not admissible. On the other hand, $\lambda:=\lambda_{\ell-2}+\lambda_{\ell}$ is not admissible by Example 3.3. While for $\lambda:=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ one notices that
@@
which says $\lambda_{\ell-3}+2\lambda_{\ell}\prec\lambda$. Since the stabilizer of $\lambda_{\ell-3}+2\lambda_{\ell}$ is of type $B_{\ell-4}\times B_{2}$, we have that
$\dim L(\lambda)\geq l(\lambda_{\ell-3}+2\lambda_{\ell})=2\ell(\ell-1)(\ell-2)(\ell-3)>\ell^{4},$
by observing $2(1-1/\ell)(1-2/\ell)(1-3/\ell)>1$ for any $\ell\geq 12$. This implies that $\lambda=\lambda_{\ell-2}+\lambda_{\ell-1}+\lambda_{\ell}$ is not admissible either.
@@
forcing then $\dim L(\lambda)\geq l(\lambda_{\ell-4})>\ell^{4}$. Therefore, the weigths in Proposition 4.4 together with the weigths in [13, Table 3] exhausts the admissible weights which conclude the proof. ∎
Finally, we will deal with the rest of the classic groups of Lie type in the next subsection.
@@ 4.2. Type $C_{\ell}$ and $D_{\ell}$
Let $G$ be a finite group of Lie type $C_{\ell}$ or $D_{\ell}$ over $\mathbb{\overline{F}}_{p}$ with $p>2$. If $\ell\geq 12$ then the following list of weights
[…]
exhausts the admissible weights for $G$.
@@
The tables in Appendix contain all the information to determine bounds for the involved multiplicities in the dimensions of the weights spaces $V(\lambda)$ and $L(\lambda)$. Hence, we may deduce that the aforementioned weights are effectively admissible. Next we shall prove that the list is exhausted respect to admissible ones. To this end, let
@@
Contrary to the previous case, here, the weights $\lambda=\lambda_{1}$, $\lambda=\lambda_{2}$ are admissible for $12\leq\ell\leq 17$, since the dimension of $L(\lambda_{i})$ is $2^{\ell-1}$, for $i=1,2$, by appealing to [9, Lemma 2.3.2]
Thereby, from the dominant weights of the form $\lambda=a_{\ell-3}\lambda_{\ell-3}+a_{\ell-2}\lambda_{\ell-2}+a_{\ell-1}\lambda_{\ell-1}+a_{\ell}\lambda_{\ell}$ on, the result can be addressed as above by using Appendix -Table 2. ∎
5. Some dimension calculations and a potential application
It is well known that an explicit formula for the dimension of $L(\lambda)$ is unknown, or at least there is no known formula holding for an arbitrary $\lambda$. Nevertheless, Theorem 3.4 shows all these dimensions when $G$ is of type $A_{\ell}$ (See also [25]). On the other hand, computing such dimensions for certain weights spaces seems to present major challenges when $G$ is not of type $A_{\ell}$, where progress has been done case-by-case according to $G$ and very special weights $\lambda$ (See, for instance, [7, 8, 9, 13, 16, 23, 24, 26, 27, 32]) However, for $p$ sufficiently large, we are able to compute some of them by extending arguments of [24, §5], where dimension formulas for the admissible weights 0, $\lambda_{\ell}$, $\lambda_{\ell-1}$ and $2\lambda_{\ell}$ have been obtained.
5.1. Type $C_{\ell}$ and $D_{\ell}$